OpenStudy (anonymous):

Coach Johnson is 10 years older than Tim. When Tim started football 3 years ago, Coach Johnson was twice his age. How old is Tim now?

OpenStudy (raden):

it can represented as : J = T+10 ... (1) J-3 = 2(T-3) ...(2)

OpenStudy (raden):

with T is Tim's old and J is Johnson's old solve for T

OpenStudy (raden):

subtitute the 1st equation to the 2nd eq J-3 = 2(T-3) T+10-3 = 2(T-3) T+7 = 2T - 6 it should helps continued ....